# ClipRover: Zero-shot Vision-Language Exploration and Target Discovery by Mobile Robots

### v1

@@ 5.3 Performance Optimization
Focusing on reducing decision delay and increasing data throughput, our proposed framework divides the navigation pipeline into four major nodes: pre-processing, inference, correlation, and decision-making. Computationally intensive tasks are offloaded to the GPU to exploit parallel processing capabilities. These optimizations were implemented on a power-limited embedded system, balancing efficiency and performance. With the proposed optimizations, we achieve: (i) decision delay: $252.10$ ms, an $900\%$ improvement; and (ii) throughput: $5.01$ FPS (frames per second), representing a $400\%$ improvement compared to the sequential CPU-based implementation. These results, detailed in Table 2, demonstrate the viability of our VLM-based navigation pipeline for real-time robotic applications under computational constraints.
Table 2: Performance margins of the navigation pipeline of ClipRover at different optimization levels. The acronyms: [ P ] Preprocess, [ I ] Inference, [ C ] Correlation, and [ D ] Decision.
@@
5.4 Usage of 2D LiDAR
The proposed ClipRover pipeline use only monocular RGB data for VLN navigation decisions. Nevertheless, we installed a 2D $360$° scanning LiDAR on the robot for experimental safety, map generation, and comparative analyses. The specific LiDAR functions are listed below.
@@ 5.4 Usage of 2D LiDAR
    Visualization of map and robot trajectory. Mapping is performed offline using recorded LiDAR and odometry data, with subsequent trajectory analysis to evaluate navigation efficiency and overall performance. We utilize the SLAM Toolbox [49] to generate maps and trajectories, implementing a two-pass process to enhance the quality of the results. In the first pass, the SLAM Toolbox operates in offline mapping mode, creating a high-resolution 2D map of the environment. In the second pass, the pre-constructed map is loaded and run in localization-only mode to generate accurate trajectories using LiDAR localization. Despite this two-pass process, the trajectories generated often exhibit zig-zag patterns caused by a mismatch between SLAM drift correction frequency and the robot’s odometry update rate. These patterns are not observed in the actual motion of the robot, still they affect our performance evaluation due to increased lengths of the projected paths. To mitigate this, we introduce a stride parameter to smooth the trajectory. This smoothing parameter reduces the projected trajectory length and better matches the robot’s actual travel distance.
@@
    Comparison with simulated algorithms. A 2D map generated using LiDAR data is employed to simulate and evaluate traditional range sensor-based map traversal and path-planning algorithms for performance comparison. The LiDAR data collected during ClipRover experiments was used to reconstruct the exploration map, which is reused for simulations. This approach ensured that all comparisons were conducted on the same map, providing a fair and consistent evaluation of the different algorithms.
5.5 Wave Front Simulation
We further design a “Wave Front” simulation package to serve as a general baseline metric for comparison between different methods. It utilizes the concept of classic wave function [51] and interprets it as probability distribution, allowing simultaneous exploration of infinity many directions driven by the following equation:
@@ 5.5 Wave Front Simulation
At time $T=0$, a probability distribution is initialized as a Gaussian distribution centered at the robot’s starting location, with a standard deviation set to half of the robot’s size; the total probability is normalized to $1$. At each iteration, the wave function is multiplied by an inverse Gaussian function centered around the target location, effectively reducing the total probability on the map and simulating a draining effect. The simulation terminates when the total remaining probability falls below a defined threshold ($1e^{-4}$ in our experiments). The outcome is a 2D probability distribution $p(t)$ indicating the likelihood of the robot reaching the target at each time step. The mean and standard deviation of this distribution are used for comparison with other trajectory-based algorithms.
At a high level, this algorithm simulates the behavior of infinitely many wall-bouncing robots moving in all possible directions. Notably, the drain operation creates a uniform gradient toward the target, subtly guiding the robot. As a result, the algorithm performs slightly better than a purely randomized wall-bouncing simulation, which lacks any target information. The travel distance when the first non-zero probability is detected is used as the baseline distance for each origin-target combination. For relative performance comparisons, as shown earlier in Fig. 2, the travel distances are normalized by this baseline distance before aggregating the results.
6 Experimental Analyses
6.1 Experiment Setup
Real-world experiments were conducted in an indoor workspace, as depicted in Fig. 10 a. This environment was selected for its cluttered layout and complex details, which include numerous obstacles, potential traps, and loops, providing a challenging setting for comprehensive analysis. As shown in Fig. 10, the robot was tasked with exploring the lab space while searching for a designated target–a toy bear approximately $20$ cm tall and $10$ cm wide, chosen for its distinctive appearance within the test scene. For each task, the source (robot’s starting position) and destination (target location) were selected from five predefined regions (e.g., SW, C, NW, NE, SE), as labeled in Fig. 10 c.
Figure 10: The environment used for real world exploration and discovery experiments: (a) An overview of the machine hall. (b) A photo of a toy bear that is used as the discovery target. (c) A 2D map of the experimental space, with major corner locations annotated in blue; the area utilized for the experiments is enclosed within the dashed red line.
6.2 Performance Evaluation
Algorithms for comparison. We compare the performance of the proposed ClipRover system with six widely used map-traversal and path-finding algorithms. For fair comparisons, a novel 2D simulation framework, RoboSim2D, is developed. It utilizes 2D LiDAR maps generated from recorded scans of the same environment used in real-world experiments. Additionally, the size of the simulated robot was configured to match the dimensions of the physical robot, ensuring consistency across simulations and the real platform.
For map traversal algorithms [56], we consider:
  1. $1$.
  2. $2$.
  3. $3$.
    Wave Front: Starts as a Gaussian probability distribution, then spreads outwards and bounces off obstacles. A detailed explanation is in Sec. 5.5.
Additionally, we compare the following Bug Algorithms [45, 47]:
  1. $1$.
  2. $2$.
  3. $3$.
Criteria of failure. Due to practical limitations, a mission is considered a failure if the distance traveled exceeds a predefined upper limit. For real-world experiments conducted in the test area (approximately $12\times 16$ m), this limit is set to $100$ meters. In simulations, the limit is set to $1,000$ meters for random walk, wall-bouncing, and wave-front algorithms. For Bug algorithms, loop detection is employed to identify endless loops, which are classified as mission failures. Since failed missions yield infinite travel distances, they are excluded from the metrics presented in Table 3. Instead, they are reported separately in the overall performance Table 4 and visualized in the relative performance plot in Fig. 2.
@@ 6.2 Performance Evaluation
       Task Setup ClipRover ★ Random Walk Wall Bounce Wave Front Bug0 Bug1 Bug2
@@
Table 4: Quantitative performance of ClipRover and other algorithms in comparison. Metrics used are defined in Eq. 4.
Method ${}^{\text{Success}}_{\text{Rate}}$ ${}^{\text{Mean Inverse}}_{\text{Path Length}}$ ${}^{\text{Additional}}_{\text{Requirements}}$
  ClipRover ★ $\phantom{0}\mathbf{95.00\%}$   $\,\,\mathbf{2.66}$±$\pm\phantom{0}1.73$        None
Random Walk $\phantom{0}50.00\%$ $10.77$±$\pm\phantom{0}6.26$
$\phantom{0}80.00\%$ $20.56$±$\pm 14.65$
[…]
$\phantom{0}80.00\%$ $18.71$±$\pm 13.66$
Wave Front $\phantom{0}50.00\%$ $13.86$±$\pm\phantom{0}5.31$
$\phantom{0}80.00\%$ $27.30$±$\pm 10.11$
Bug0 $\phantom{0}45.00\%$   $\,\,\mathbf{1.14}$±$\pm\phantom{0}0.08$ Target location
[…]
$\text{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{l_{i}}{\max(p_{i},l_{i})}.$ (2)
Note that, for each source-target combination, the first contact distance of WaveFront simulation is used as the baseline distance $D_{\text{baseline}}$. With the choice of $l=D_{\text{baseline}}$, we have $p_{i}>l,~{}\forall p_{i}\in\{p\}$; hence, the above equation is simplified to:
[…]
As a consequence, the SPL metric equates to an inverse proportional relation between two quantities, namely the success rate $R$ and the mean inverse path length $\overline{L}$, as follows:
$\overline{L}=\frac{\text{SPL}}{R}.$ (5)
However, experimental data reveal that the curvature of the SPL function does not align well with the observed $R$$\overline{L}$ curve shown in Fig. 2. To better model the observed data, we introduce a new metric $\bm{\varepsilon}$ to represent the amount of information available to a given algorithm. It has a range of $[0,1]$. When $\bm{\varepsilon}=0$, the algorithm operates without environmental information and lacks sensing capabilities beyond collision detection, as seen in methods like random walk and wall bounce. As $\bm{\varepsilon}$ approaches $1$, it gains more information or has higher environment sensing capability, and the performance is expected to increase accordingly. Considering the impact of $\bm{\varepsilon}$, the revised relationship reduces to:
$\overline{L}=\frac{\bm{\varepsilon}^{\prime}}{(kR+t)^{\beta}}$ (6)
[…]
where $k$, $t$, and $\alpha$ are hyper-parameters specific to the environment. They are derived using linear regression with $\bm{\varepsilon}=0$; additionally, $\omega$ is solvable by setting $\bm{\varepsilon}=\overline{L}=R=1$.
Figure 11: Sample results of autonomous exploration and target discovery by ClipRover compared to Bug algorithms. The trajectories of each algorithm are overlaid on a 2D map of the test environment. Circled numbers indicate the look around operations of ClipRover and the corresponding waypoints sequence it followed. For Bug algorithms, the red, green, and blue trajectories represent paths traversed by Bug0, Bug1, and Bug2, respectively. Best viewed digitally at 2×2$\times$2 × zoom for clarity; more results are in the Appendix. A video demonstration is provided here: https://youtu.be/84XLM-GbhS8.
6.3 Qualitative and Quantitative Analyses
We conduct extensive real-world experiments with over $60$ trials based on various combinations of source-target locations shown in Fig. 10. For the proposed ClipRover system, three trials were performed for each origin-target pair. In comparison, $200$ trials were recorded for each random walk experiment, while $180$ evenly distributed headings were recorded for each combination of the wall bounce runs.
Samples of the trajectories traversed by ClipRover and other algorithms are presented in Fig. 11, overlaid on a 2D map of the environment. Under normal navigation mode, ClipRover exhibits a human-like motion strategy by navigating near the center of open spaces, enhancing exploration efficiency, and reducing the likelihood of becoming trapped by obstacles. In contrast, path-finding algorithms often follow the contours of obstacles due to their lack of semantic scene understanding. During look-around operations (see Fig. 8 and Fig. 9), ClipRover accurately discriminates between navigable spaces and obstacles. Additionally, it demonstrates a preference for unexplored (unfamiliar) areas over previously visited (familiar) regions, contributing to improved efficiency compared to traditional map-traversal algorithms.
Out of all $60$ trials, 3 failure cases were recorded, yielding $95\%$ overall success rate for ClipRover. Among them, two were caused by the trajectory length exceeding the upper limit, while the other one was caused by the robot getting jammed by a reflective metal tank, which was not detected by either the VLN pipeline or the LiDAR proximity switch.
The quantitative results categorized by source and target locations are presented in Table 3. To ensure comparability across different source-target pairs, the travel distance for each row is normalized by the baseline travel distance and then aggregated into Table 4. For randomized algorithms, the success rate $R$ is a configurable hyperparameter. Sample results fo $R=50\%$ and $R=80\%$ are provided to represent their underlying performance in Table 4, with the corresponding equipotential curves previously shown in Fig. 2. In contrast, deterministic algorithms (such as Bug variants) have a fixed success rate. Their performances are reported directly in Table 4 and visualized in Fig. 2 for comparison.
As these results demonstrate, ClipRover averages $2.66\times D_{\text{baseline}}$ travel distance before reaching the target, while the next best score is from Wall Bounce, with $9.58\times D_{\text{baseline}}$ at a significantly lower success rate ($50\%$). Note that, it achieves this performance gains despite having the same prior information as the map traversal algorithms. On the other hand, although the Bug* algorithms uses additional information about the environment map and target location, ClipRover ($95\%$) still outperforms outperforms Bug1 ($6.34\times D_{\text{baseline}}$) by trajectory lengths, and offers significantly higher success rates than Bug0 ($45\%$) and Bug2 ($80\%$) algorithms.
In summary, the proposed system outperforms traditional map traversal algorithms by significant margins in both success rate and trajectory efficiency, given the same amount of information. Despite the inherent disadvantage of operating without prior target knowledge or precise localization, the system achieves performance comparable to path-finding algorithms. Notably, in many scenarios, it surpasses path-finding algorithms in either trajectory efficiency or success rates. These results highlight the effectiveness of the proposed pipeline for simultaneous exploration and target discovery tasks.
Table 5: A qualitative comparison of features and prerequisites among state-of-the-art VLN systems.
CLIP on Wheels [17] Object Unspecified RGB-D Camera ${}^{\text{Revalance map of}}_{\text{the current View}}$ Simulation ${}^{\text{Accuracy \&}}_{\text{Path length}}$
[…]
6.4 Comparative Analyses of SOTA VLNs
We also compare ClipRover with state-of-the-art VLN systems based on their features and prerequisites. The characteristics of each system are summarized in Table 5. Compared to existing systems, ClipRover uniquely requires zero prior knowledge of the task space and operates solely with monocular RGB perception. Additionally, contemporary VLN systems do not directly produce low-level motion commands but instead generate high-level instructions that depend on another traditional localization and navigation pipeline for the eventual motion execution. In contrast, ClipRover directly outputs low-level motion commands, eliminating the need for auxiliary control and localization systems. This feature makes ClipRover particularly suitable for rapid deployment on low-cost robots with limited computational and sensing resources.
[…]
7 Limitations, Failure Cases, and Potential Improvements of ClipRover
7.1 Texture-less or Transparent Surfaces
As a vision-driven system, the proposed ClipRover pipeline encounters challenges in environments with texture-less or transparent surfaces. When presented with a blank or non-informative image, the inherent VLM struggles to generate meaningful responses, disrupting the entire pipeline. For instance, as demonstrated in Fig. 9 d (C1), the robot mistakenly classified a water tank with transparent walls as navigable space. This error occurred because the VLM perceived the interior of the tank, visible through its transparent walls, as accessible terrain. To mitigate this issue, a standard deviation threshold was introduced to filter out ambiguous tiles. While this approach improved performance, the system still exhibits a general tendency to misclassify transparent or uniformly colored surfaces, such as blank walls or white floors, as navigable space. These challenges persist, particularly when the robot is positioned close to such surfaces, where the lack of texture further confuses the VLM. Note that this issue with transparent and texture-less walls has been an open problem for 2D indoor navigation by mobile robots [78].
7.2 Open Space and Artificially Constructed Mazes
Moreover, our experiments in artificially constructed mazes, such as those built from paper boxes, revealed limitations in the VLM’s ability to differentiate between the maze walls such as monochrome cardboards, and floors. As a result, the robot either failed to navigate or attempted to collide with the maze walls. This behavior can be attributed to the lack of similar examples in the VLM’s training data, which likely did not include artificially constructed environments of this nature. Consequently, the system is better suited to real-world scenarios where the VLM has been trained on relevant, diverse visual data. This observation suggests that our proposed system will work most effectively in real-world environments where VLMs generally work well.
7.3 Familiarity Saturation
7.4 Adaptive Slicing Strategies
In our current implementation of ClipRover, a fixed slicing strategy is used based on the camera frame’s aspect ratio, which generally performs well in general. However, in specific scenarios involving narrow navigable spaces (e.g., a $50$ cm gap between two obstacles), the robot often fails to recognize these spaces as traversable. This is because the narrow passage does not occupy an entire tile, leaving obstacles visible in all tiles and leading the system to avoid the space rather than navigate through it. Future improvements could include the development of advanced slicing strategies that dynamically adapt the region of interest (ROI) based on the surrounding environment. This could be achieved using sparse object detection (bounding box) or pixel-wise segmentation models, enabling better handling of narrow spaces.
7.5 High-Level Reasoning and Task Narration
One of the most promising extensions of this work lies in incorporating high-level reasoning and task narration using an integrated LLM. With this addition, the robot could leverage common-sense reasoning to make more intelligent navigation decisions. For example, when prompted to locate an item in a different room, the robot could infer that heading toward a door is the most logical action. Moreover, in bandwidth-limited applications, the LLM could provide real-time task narration by interpreting visual embeddings to describe mission progress. This approach would significantly reduce the communication bandwidth required compared to streaming raw video. While the architecture lays the foundation for these enhancements, their implementation remains a future work. By addressing these limitations and pursuing the outlined improvements, ClipRover can evolve into a more robust, efficient, and versatile platform for VLN in complex real-world environments.
8 Conclusion
In this paper, we introduced ClipRover, a novel navigation pipeline designed for simultaneous exploration and target discovery by autonomous ground robots in unknown environments, leveraging the power of VLMs. Unlike traditional approaches that decouple exploration from path planning and rely on inefficient algorithms, ClipRover enables zero-shot inference and efficient navigation using only monocular vision, without prior maps or target-specific information. To validate our approach, we developed a functional prototype UGV platform named Rover Master, optimized for general-purpose VLN tasks in real-world environments. Extensive evaluations demonstrated that ClipRover outperforms state-of-the-art map traversal algorithms in efficiency and achieves comparable performance to path-planning methods that rely on prior knowledge. The integration of a CLIP-based VLM into a real-time navigation system underscores the potential of ClipRover to advance intelligent robotic exploration. As a modular and flexible framework, ClipRover lays the groundwork for future research in applying VLMs to more complex and dynamic robotic applications such as autonomous warehousing, security patrolling, and smart home assistance.
Acknowledgments
This work is supported in part by the National Science Foundation (NSF) grants #$2330416$, #$19244$; the Office of Naval Research (ONR) grants #N000142312429, #N000142312363; and the University of Florida (UF) ROSF research grant #$132763$.
Appendix
Some additional results are shown in the following page. Please refer to Fig. 11 for the relevant discussions. A video demonstration is provided here: https://youtu.be/84XLM-GbhS8.
Figure 12: Qualitative comparison of all $20$ trajectories for different source-target combinations among five points (C, N.W., N.E., S.W., S.E.) is illustrated (continued on the next page).
Figure 13: Qualitative comparison of all $20$ trajectories for different source-target combinations among five points (C, N.W., N.E., S.W., S.E.) is illustrated. See Table 3 for a quantitative comparison.

### v2

@@ Appendix C Performance Optimization
Focusing on reducing decision delay and increasing data throughput, our proposed framework divides the navigation pipeline into four major nodes: pre-processing, inference, correlation, and decision-making. Computationally intensive tasks are offloaded to the GPU to exploit parallel processing capabilities. These optimizations were implemented on a power-limited embedded system, balancing efficiency and performance. With the proposed optimizations, we achieve: (i) decision delay: $252.10$ ms, a $900\%$ improvement; and (ii) throughput: $5.01$ FPS (frames per second), representing a $400\%$ improvement compared to the sequential CPU-based implementation. These results, detailed in Table 4, demonstrate the viability of our VLM-based navigation pipeline for real-time robotic applications under computational constraints.
Table 4: Performance margins of the navigation pipeline of VL-Explore  at different optimization levels. The acronyms: [ P ] Preprocess, [ I ] Inference, [ C ] Correlation, and [ D ] Decision.
@@
Appendix D Usage of 2D LiDAR
The proposed VL-Explore  pipeline uses only monocular RGB data for VLN navigation decisions. Nevertheless, we installed a 2D $360$° scanning LiDAR on the robot for experimental safety, map generation, and comparative analyses. The specific LiDAR functions are listed below.
@@ Appendix D Usage of 2D LiDAR
    Visualization of map and robot trajectory. Mapping is performed offline using recorded LiDAR and odometry data, with subsequent trajectory analysis to evaluate navigation efficiency and overall performance. We utilize the SLAM Toolbox [31] to generate maps and trajectories, implementing a two-pass process to enhance the quality of the results. In the first pass, the SLAM Toolbox operates in offline mapping mode, creating a high-resolution 2D map of the environment. In the second pass, the pre-constructed map is loaded and run in localization-only mode to generate accurate trajectories using LiDAR localization. Despite this two-pass process, the trajectories generated often exhibit zig-zag patterns caused by a mismatch between SLAM drift correction frequency and the robot’s odometry update rate. These patterns are not observed in the actual motion of the robot; still, they affect our performance evaluation due to increased lengths of the projected paths. To mitigate this, we introduce a stride parameter to smooth the trajectory. This smoothing parameter reduces the projected trajectory length and better matches the robot’s actual travel distance.
@@
    Comparison with simulated algorithms. A 2D map generated using LiDAR data is employed to simulate and evaluate traditional range sensor-based map traversal and path-planning algorithms for performance comparison. The LiDAR data collected during VL-Explore  experiments was used to reconstruct the exploration map, which is reused for simulations. This approach ensured that all comparisons were conducted on the same map, providing a fair and consistent evaluation of the different algorithms.
Appendix E Simulation Algorithms
For map traversal algorithms [34], we consider:
  1. $1$.
  2. $2$.
  3. $3$.
    Wave Front: Starts as a Gaussian probability distribution, then spreads outwards and bounces off obstacles. A detailed explanation is provided below.
Additionally, we compare the following Bug Algorithms [28, 29]:
  1. $1$.
  2. $2$.
  3. $3$.
E.1 Wave Front Simulation
The “Wave Front” simulation is a custom-developed package to serve as a general baseline metric for comparison between different methods. It utilizes the concept of classic wave function [32] and interprets it as a probability distribution, allowing simultaneous exploration of infinitely many directions driven by the following equation:
@@ E.1 Wave Front Simulation
At time $T=0$, a probability distribution is initialized as a Gaussian distribution centered at the robot’s starting location, with a standard deviation set to half of the robot’s size; the total probability is normalized to $1$. At each iteration, the wave function is multiplied by an inverse Gaussian function centered around the target location, effectively reducing the total probability on the map and simulating a draining effect. The simulation terminates when the total remaining probability falls below a defined threshold ($1e^{-4}$ in our experiments). The outcome is a 2D probability distribution $p(t)$ indicating the likelihood of the robot reaching the target at a given time point. The mean and standard deviation of this distribution are used for comparison with other trajectory-based algorithms.
At a high level, this algorithm simulates the behavior of infinitely many wall-bouncing robots moving in all possible directions. Notably, the drain operation creates a uniform gradient toward the target, subtly guiding the robot. As a result, the algorithm performs slightly better than a purely randomized wall-bouncing simulation, which lacks any target information. The travel distance when the first non-zero probability is detected is used as the baseline distance for each origin-target combination. For relative performance comparisons, the travel distances are normalized by this baseline distance before aggregating the results.
Appendix F Additional Experimental Results
The full set of trajectories recorded in the experiment is shown in Fig. 14 and Fig. 15. Metrics measured for each source-target combination are shown in Table 5.
Figure 14: Trajectories of all results reported in the experiment section (part.1)
Figure 15: Trajectories of all results reported in the experiment section (part.2)
@@ Appendix F Additional Experimental Results
       Task Setup VL-Explore ★ Random Walk Wall Bounce Wave Front Bug0 Bug1 Bug2
@@
       Office Scene ($13m\times 12.5m$)
  N.W. S. E. $22.8$±$\pm\phantom{0}3.6$ $173.6$±$\pm 15.7$ $148.0$±$\pm 10.6$ $48.0$±$\pm 22.2$$\times$ $\times$ $88.4$ $64.1$ $25.4$ $34.9$
N. E. S.W. $36.3$±$\pm\phantom{0}5.7$ $243.3$±$\pm 19.8$ $207.7$±$\pm 13.5$ $56.9$±$\pm 28.2$$\times$ $17.0$ $66.6$ $88.0$ $23.9$ $35.0$
S.W. N. E. $16.9$±$\pm\phantom{0}4.3$ $263.9$±$\pm 24.4$ $268.3$±$\pm 16.9$ $65.7$±$\pm 34.3$$\times$ $14.5$ $39.2$ $47.0$ $36.0$ $31.4$
S. E. N.W. $14.5$±$\pm\phantom{0}0.7$ $266.3$±$\pm 19.2$ $290.4$±$\pm 18.5$ $72.5$±$\pm 38.6$ $15.9$$\times$ $47.4$ $40.1$ $37.7$ $25.9$